### v1

# IMPaCT: Interval MDP Parallel Construction for Controller Synthesis of Large-Scale STochastic Systems

Abstract.
This paper is concerned with developing a software tool, called IMPaCT, for the parallelized verification and controller synthesis of large-scale stochastic systems using interval Markov chains (IMCs) and interval Markov decision processes (IMDPs), respectively. The tool serves to (i) construct IMCs/IMDPs as finite abstractions of underlying original systems, and (ii) leverage interval iteration algorithms for formal verification and controller synthesis over infinite-horizon properties, including safety, reachability, and reach-avoid, while offering convergence guarantees. IMPaCT is developed in C++ and designed using AdaptiveCpp, an independent open-source implementation of SYCL, for adaptive parallelism over CPUs and GPUs of all hardware vendors, including Intel and NVIDIA. IMPaCT stands as the first software tool for the parallel construction of IMCs/IMDPs, empowered with the capability to leverage high-performance computing platforms and cloud computing services. Specifically, parallelism offered by IMPaCT effectively addresses the challenges arising from the state-explosion problem inherent in discretization-based techniques applied to large-scale stochastic systems. We benchmark IMPaCT on several physical case studies, adopted from the ARCH tool competition for stochastic models, including a 222-dimensional robot, a 333-dimensional autonomous vehicle, a 555-dimensional room temperature system, and a 777-dimensional building automation system. To show the scalability of our tool, we also employ IMPaCT for the formal analysis of a 141414-dimensional case study.
Keywords: Interval Markov chain, interval Markov decision process, automated controller synthesis, large-scale stochastic systems, parallel construction, cloud computing
1. Introduction
Large-scale stochastic systems serve as a crucial modeling framework for characterizing a wide array of real-world safety-critical systems, encompassing domains such as power grids, autonomous vehicles, communication networks, smart buildings, energy systems, and so on. The intended behavior of such complex systems can be formally expressed using high-level logic specifications, typically formulated as linear temporal logic (LTL) expressions [BK08]. Automating the formal verification and controller synthesis for these complex systems that fulfill LTL specifications is an exceedingly formidable challenge (if not impossible), primarily due to uncountable nature of states and actions in continuous spaces.
To tackle the computational complexity challenges that arise, one promising solution is to approximate original (a.k.a. concrete) systems with simpler models featuring finite state sets, commonly referred to as finite abstractions [Tab09, BYG17]. When the underlying models are stochastic, these simplified representations commonly adopt the structure of Markov decision processes (MDPs), where discrete states mirror sets of continuous states in the concrete model (similarly for inputs). In practical implementation, constructing such finite abstractions usually entails partitioning the state and input sets of concrete models according to predefined discretization parameters, as discussed in various works including  [APLS08, JP09, ZMEM+14, TMKA13, HHHK13, LZ22, Nej23, NSZ21, LZ19, LF22, NZ20, Lav19, LSAZ22].
Within the finite MDP scheme, one can (i) initially leverage it as an appropriate substitute for the original system, (ii) proceed to synthesize controllers for the abstract system, and (iii) ultimately refine the controller back over the concrete model, facilitated by an interface map. Given that the disparity between the output of the original system and that of its abstraction is accurately quantified, it becomes feasible to ensure that the concrete system fulfills the same specification as the abstract counterpart under some quantified accuracy level. However, this accuracy level is only acceptable for finite-horizon specifications since it converges to infinity as time approaches infinity (cf. (2.4)), rendering MDPs impractical for infinite-horizon properties.
As a promising alternative, interval Markov decision processes (IMDPs) [SVA06, GLD00] have emerged in the literature as a potential solution for formal verification and controller synthesis of stochastic systems fulfilling infinite-horizon specifications. Specifically, IMDPs offer a comprehensive approach by incorporating both upper and lower bounds on the transition probabilities among finite abstraction cells. This is accomplished by solving a (multi-player) game which allows to extend satisfaction guarantees to infinite time horizons. However, the construction of IMDPs is more complicated compared to traditional MDPs as it necessitates the computation of both lower and upper bounds for transition probabilities among partition cells.
Abstraction-based techniques, including those used for constructing either MDPs or IMDPs, encounter a significant challenge known as the curse of dimensionality. This phenomenon refers to the exponential growth in computational complexity as the number of state dimensions increases. To alleviate this, we offer scalable parallel algorithms and efficient data structures designed for constructing IMCs/IMDPs and automating the process of verification and controller synthesis over infinite time horizons. In particular, by dividing the computations into smaller concurrent operations, we effectively mitigate the overall complexity by a factor equivalent to the number of threads available. This approach not only enhances computational efficiency of IMC/IMDP construction, but also facilitates the practical application of these techniques in real-world scenarios with high-dimensional spaces.
1.1. Original Contributions
The primary contributions and noteworthy aspects of our tool paper include:
  1. (i)
    We propose the first tool that constructs IMCs/IMDPs for large-scale discrete-time stochastic systems while providing convergence guarantees. Our tool leverages the constructed IMCs/IMDPs for formal verification and controller synthesis ensuring the fulfillment of desired infinite-horizon temporal logic specifications, encompassing safety, reachability, and reach-while-avoid properties.
  2. (ii)
    IMPaCT is implemented in modern ISO C++ and runs in parallel using AdaptiveCpp [^1] https://github.com/AdaptiveCpp/AdaptiveCpp/ based on SYCL [^2] https://www.khronos.org/sycl/. AdaptiveCpp eliminates from the user the need to implement cross-platform flexibility manually, serving as a strong foundation for CPU and GPU implementations.
  3. (iii)
    IMPaCT leverages interval iteration algorithms to provide convergence guarantees to an optimal controller in scenarios with infinite time horizons.
  4. (iv)
    IMPaCT accepts bounded disturbances and natively supports additive noises with different practical distributions including normal and user-defined distributions.
  5. (v)
    We leverage IMPaCT across diverse real-world applications such as autonomous vehicles, room temperature systems, and building automation systems. This broadens the scope of formal method techniques to encompass safety-critical applications that require satisfaction within infinite time horizons. The outcomes demonstrate significant efficiency in computational time.
The source code for IMPaCT along with comprehensive instructions on how to build and operate it can be located at:
https://github.com/Kiguli/IMPaCT
1.2. Related Literature
IMDPs have become the source of significant interest inside the formal methods community over the past few years. In particular, IMDPs have been used for model-based verification and control problems [CHK13, DC20, HHS+16, WMK19, DLML23] as well as more recently being used for data-driven learning problems [JZC22, LSFZ22, RAM23]. There exists a limited set of software tools available for the formal verification and controller synthesis of stochastic systems, encompassing various classes of stochastic models, using abstraction-based techniques. FAUST2 [SGA15] constructs finite MDPs for continuous-space discrete-time stochastic processes and conducts formal analysis for safety and reachability specifications. Nonetheless, the original MATLAB implementation of FAUST2 encounters scalability issues, particularly for large models, due to the curse of dimensionality.
StocHy [CA19] offers formal verification and synthesis frameworks for discrete-time stochastic hybrid systems via finite MDPs. While a small segment of StocHy addresses IMDPs, its primary implementation relies on the value iteration algorithm [LAB15], which lacks convergence guarantees when dealing with infinite-horizon specifications. This limitation has been widely pointed out in the relevant literature (see e.g., [HM14, BKL+17, HM18]). For the assurance of convergence to an optimal controller, it is essential to utilize interval iteration algorithms. This is a key feature of our tool on top of offering parallelization, which sets IMPaCT apart from StocHy, where value iteration is the default choice without any general parallelization capability. In particular, IMPaCT offers a notably more comprehensive and versatile approach for addressing infinite-horizon specifications and stands as the first tool dedicated to IMC/IMDP construction with convergence guarantees.
A recent update to PRISM [KNP11] supports robust verification of uncertain models including IMDPs. However, PRISM requires the states and transitions of the IMDP to be described explicitly, which is cumbersome when the state and input spaces are large. Furthermore, our tool introduces parallel implementations of IMDP abstraction and synthesis algorithms, efficiently automating the IMDP construction process. It is noteworthy to mention that AMYTISS [LKSZ20] also employs parallel implementations but exclusively for constructing MDPs using PFaces [KZ19], built upon OpenCL, without any support for IMDP models. In addition, our tool is developed using SYCL, which facilitates parallel processing and offers broader utility than PFaces. In particular, AdaptiveCpp (formerly OpenSYCL) [AH20, AH23], employed in IMPaCT, provides a sturdy foundation for flexible cross-platform parallel processing on CPUs, GPUs and hardware accelerators (HWAs) from all major hardware vendors such as Intel, NVIDIA, etc., which is not the case in AMYTISS.
There exist a few more software tools for formal verification and controller synthesis of stochastic systems within diverse model classes, all without employing IMDP approaches. In this regard, SReachTools [VGO19] conducts stochastic reachability analysis specifically for linear (time-varying) discrete-time stochastic systems. ProbReach [SZ15] is developed for the probabilistic reachability verification of stochastic hybrid systems. SReach [WZK+15] addresses probabilistic bounded reachability problems specifically within nonlinear hybrid automata affected by parametric uncertainty. Modest Toolset [HH14] enables modeling and analysis of distributed stochastic hybrid systems. The last three ARCH workshops focused on friendly tool competitions over stochastic models are reported in [ABC+, ABD+22, ABC+23].
1.3. Paper Organization
This paper is structured as follows. In Section 2, we define discrete-time stochastic control systems and their equivalent representations as continuous-space MDPs. In Section 3, we define IMDPs, while presenting a serial algorithm for IMDP construction. Section 4 is dedicated to the parallel algorithms of IMDP construction. In Section 5, we discuss serial algorithms for controller synthesis via constructed IMDPs, enforcing safety, reachability, and reach-while-avoid properties. Section 6 offers the parallel synthesis algorithms using constructed IMDPs. Section 7 elaborates on the process of loading and saving respective files throughout various stages of the abstraction and synthesis procedure. Section 8 showcases the outcomes of IMPaCT when implemented on well-known case studies and benchmark scenarios.
1.4. Preliminaries
[…]
2. Discrete-Time Stochastic Control Systems
A formal description of discrete-time stochastic control systems, serving as the underlying dynamics of our tool, is presented in the following definition.
Definition 2.1 (dt-SCS).
A discrete-time stochastic control system (dt-SCS) is a quintuple
$\Sigma=(X,U,W,\varsigma,f),$ (2.1)
where,
  • $X\subseteq\mathbb{R}^{n}$ is a Borel space as the state set, with $(X,\mathscr{B}(X))$ being its measurable space;
  • $U\subseteq\mathbb{R}^{m}$ is a Borel space as the input set;
  • $W\subseteq\mathbb{R}^{p}$ is a Borel space as the disturbance set;
  • $\varsigma$ is a sequence of independent and identically distributed (i.i.d.) random variables from a sample space $\Omega$ to a measurable set $\mathcal{V}_{\varsigma}$
    $\varsigma:=\{\varsigma(k)\!:\Omega\rightarrow\mathcal{V}_{\varsigma},~k\in\mathbb{N}\};$
  • $f\!:X\times U\times W\times\mathcal{V}_{\varsigma}\rightarrow X$ is a measurable function characterizing the state evolution of the system.
For a given initial state $x(0)\in X$, an input sequence $u(\cdot):\Omega\rightarrow U$, and a disturbance sequence $w(\cdot):\Omega\rightarrow W$, the state evolution of $\Sigma$ is characterized by
$\Sigma\!:x(k+1)=f(x(k),u(k),w(k),\varsigma(k)),\quad k\in\mathbb{N}.$ (2.2)
To facilitate a more straightforward presentation of our contribution, we illustrate our algorithms by incorporating stochasticity with normal distributions. Nevertheless, our tool is capable of handling problems involving any arbitrary distributions via a custom user-defined distribution.
A dt-SCS has been shown to be equivalent to a continuous-space Markov decision process [Kal21] as the following definition, where a conditional stochastic kernel $T$ captures the evolution of $\Sigma$ and can be uniquely determined by the pair $(\varsigma,f)$ from (2.1).
Definition 2.2 (Continuous-Space MDPs).
A continuous-space Markov decision process (MDP) is a quadtuple
$\Sigma=(X,U,W,T),$ (2.3)
where,
  • $X$, $U$, and $W$ are as in Definition 2.1;
  • $T\!:\mathscr{B}(X)\times X\times U\times W\rightarrow[0,1]$ is a conditional stochastic kernel that assigns any $x\in X$, $u\in U$, and $w\in W$, a probability measure $T(\cdot|x,u,w)$, on the measurable space $(X,\mathscr{B}(X))$ so that for any set $A\in\mathscr{B}(X)$,
    $\mathbb{P}\Big{\{}x(k+1)\in A\,\big{|}\,x(k),u(k),w(k)\Big{\}}=\int_{A}T(\text{d}x(k+1)\,\big{|}\,x(k),u(k),w(k)).$
To construct a traditional finite MDP (i.e., an MDP with finite spaces), the continuous spaces $X,U,W$ are constructed from a finite number of partitions $\textbf{X}^{i},\textbf{U}^{i},\textbf{W}^{i}$, where $X=\cup_{i=0}^{n_{u}}\textbf{X}^{i},U=\cup_{i=0}^{n_{x}}\textbf{U}^{i},W=\cup_{i=0}^{n_{w}}\textbf{W}^{i}$. The number of partitions $n_{x},n_{u},n_{w}$ can be computed based on discretization parameters $\eta_{x},\eta_{u},\eta_{w}$ which defines the size of regions $\textbf{X}^{i},\textbf{U}^{i},\textbf{W}^{i}$, respectively. Each finite partition can be identified by some representative points $\hat{x}_{i}\in\textbf{X}^{i}$, $\hat{u}_{i}\in\textbf{U}^{i}$, $\hat{w}_{i}\in\textbf{W}^{i}$; the collections of all representative points can now be considered as the finite sets $\hat{X},\hat{U},\hat{W}$ of the finite MDP. Using a map $\Xi:X\rightarrow 2^{X}$, one can assign any continuous state $x\in X$ to the partition $\textbf{X}^{i}$, where $x\in\textbf{X}^{i}$. Additionally, a map $\Pi_{x}:X\rightarrow\hat{X}$ assigns any continuous state $x\in X$ to the representative point $\hat{x}\in\hat{X}$ of the corresponding partition containing $x$. The map $\Pi_{x}$ satisfies the inequality
$\lVert\Pi_{x}(x)-x\rVert_{2}\leq\eta_{x},\quad\quad\forall x\in X,$
with $\eta_{x}$ being a state discretization parameter. Using these mappings, the transition function $\hat{f}$ and the transition probability matrix $\hat{T}$ within the finite MDP construction are defined as $\hat{f}(\hat{x},\hat{u},\hat{w},\varsigma)=\Pi_{x}(f(\hat{x},\hat{u},\hat{w},\varsigma))$ and $\hat{T}(\hat{x}^{\prime}|\hat{x},\hat{u},\hat{w})=T(\Xi(\hat{x}^{\prime})|\hat{x},\hat{u},\hat{w})$, respectively [LSZ18, Alg. 1 & Thm. 2.2].
Of particular importance to this work, an accuracy level $\rho$ regards the difference between probabilities of satisfaction of the desired specification $\psi$ over the continuous-space system $\Sigma$ and its finite MDP counterpart $\hat{\Sigma}$. This accuracy level $\rho$ can be computed a-priori as the product of the Lipschitz constant $\mathcal{H}$ of the stochastic kernel, the Lebesgue measure $\mathscr{L}$ of the state space, the state discretization parameter $\eta_{x}$, and the finite horizon $\mathcal{K}$, as the following:
$\big{|}\mathbb{P}(\Sigma\models\psi)-\mathbb{P}(\hat{\Sigma}\models\psi)\big{|}\leq\rho=\mathcal{K}\mathcal{H}\mathscr{L}\eta_{x}.$ (2.4)
3. Interval Markov Decision Processes
The finite MDP presented in Section 2 is not suitable for infinite-horizon problems due to the inclusion of the time horizon $\mathcal{K}$ in (2.4). In particular, as $\mathcal{K}\rightarrow\infty$ then $\rho\rightarrow\infty$, the distance between probabilities of satisfaction over concrete system $\Sigma$ and its finite MDP $\hat{\Sigma}$ converges to infinity, implying that the abstraction will be of no use to providing satisfaction guarantees. To alleviate this, a continuous-space MDP $\Sigma$ in (2.3) can be finitely abstracted by an interval Markov decision process. Specifically, IMDPs provide bounds over the transition probability of the stochastic kernel $T$, offering a reliable model for analyzing infinite-horizon specifications, as outlined in the subsequent definition.
Definition 3.1 (IMDPs).
An interval Markov decision process (IMDP) is defined as a quintuple
$\hat{\Sigma}=(\hat{X},\hat{U},\hat{W},\hat{T}_{{\min}},\hat{T}_{{\max}}),$
where,
  • $\hat{X}=\{\hat{x}_{0},\hat{x}_{1},\ldots\hat{x}_{n_{x}}\}$, with $\hat{x}_{i}$ being the representative point within $\textbf{X}^{i}$, where $X=\cup_{i=0}^{n_{x}}\textbf{X}^{i}$;
  • $\hat{U}=\{\hat{u}_{0},\hat{u}_{1},\ldots\hat{u}_{n_{u}}\}$, with $\hat{u}_{i}$ being the representative point within $\textbf{U}^{i}$, where $U=\cup_{i=0}^{n_{u}}\textbf{U}^{i}$;
  • $\hat{W}=\{\hat{w}_{0},\hat{w}_{1},\ldots\hat{w}_{n_{w}}\}$, with $\hat{w}_{i}$ being the representative point within $\textbf{W}^{i}$, where $W=\cup_{i=0}^{n_{w}}\textbf{W}^{i}$;
  • $\hat{T}_{{\min}}$ is a conditional stochastic kernel for the minimal transition probability, computed as
    $\hat{T}_{{\min}}(\hat{x}^{\prime}\,|\,\hat{x},\hat{u},\hat{w})=\min\limits_{x\in\Xi(\hat{x})}~T(\Xi(\hat{x}^{\prime})\,|\,x,\hat{u},\hat{w}),\quad\forall\hat{x},\hat{x}^{\prime}\in\hat{X},~\forall\hat{u}\in\hat{U},~\forall\hat{w}\in\hat{W},$
    with $x\in\Xi(\hat{x})$, and where the map $\Xi:X\rightarrow 2^{X}$ assigns to any $x\in X$, the corresponding partition element it belongs to, i.e., $\Xi(x)=\textbf{X}^{i}$ if $x\in\textbf{X}^{i}$;
  • $\hat{T}_{{\max}}$ is a conditional stochastic kernel for the maximal transition probability, computed similarly as
    $\hat{T}_{{\max}}(\hat{x}^{\prime}\,|\,\hat{x},\hat{u},\hat{w})=\max\limits_{x\in\Xi(\hat{x})}~T(\Xi(\hat{x}^{\prime})\,|\,x,\hat{u},\hat{w}),\quad\forall\hat{x},\hat{x}^{\prime}\in\hat{X},~\forall\hat{u}\in\hat{U},~\forall\hat{w}\in\hat{W},$
    $\hat{T}_{{\min}}\leq\hat{T}_{{\max}},~\text{and}~\sum\limits_{\hat{x}^{\prime}\in\hat{X}}\hat{T}_{{\min}}(\hat{x}^{\prime}\,|\,\hat{x},\hat{u},\hat{w})\leq 1\leq\sum\limits_{\hat{x}^{\prime}\in\hat{X}}\hat{T}_{{\max}}(\hat{x}^{\prime}\,|\,\hat{x},\hat{u},\hat{w}).$
Although IMDPs incur higher computational costs compared to traditional MDPs, leveraging lower and upper bound probabilities for state transitions facilitates the resolution of infinite-horizon control problems. It is worth mentioning that since $\hat{X},\hat{U},\hat{W}$ are all finite sets, $\hat{T}_{{\min}}$ and $\hat{T}_{{\max}}$ can be represented by static matrices of size $(n_{x}\times n_{u}\times n_{w})$ by $n_{x}$. This enables the use of powerful iterative algorithms.
3.1. Temporal Logic Specifications
Here, we define the specifications of interest handled by IMPaCT. The desired specification $\psi$ is codified using linear temporal logic (LTL) [BK08]. Of particular interest are the properties described via logical operators including always $\square$, eventually $\lozenge$, and until $\mathbin{\sf U}$.
Definition 3.2 (Specifications).
The specifications of interest $\psi$, handled by IMPaCT, are defined as
  • $\psi:=\square{\mathcal{S}}$ safety; the system should always remain within a safe region ${\mathcal{S}}\subseteq X$.
  • $\psi:=\lozenge\mathcal{T}$ reachability; the system should eventually reach some target region $\mathcal{T}\subseteq X$.
[…]
When constructing IMCs/IMDPs, we aim at labeling the states within the state space based on the specification, especially beneficial for later-stage verification or controller synthesis processes. To do so, with a slight abuse of notation, we consider states $\hat{x}\in\mathcal{S}$ to refer explicitly to states in the state space considered to be safe, i.e., for safety specifications these states belong to the safe region $\mathcal{S}$. We now relabel states from the state space that belong to the target set as target states $\hat{r}\in\mathcal{T}$ and assume that all of these states are absorbing states, i.e., if the system reaches the target region, it will remain within that region. Similarly, we relabel any states from the state space that also belong to the avoid region as avoid states $\hat{a}\in\mathcal{A}$, treating the avoid region as an absorbing area. Due to the absorbing properties, the avoid and target regions are commonly modeled as a single state to simplify the algorithms. Transition probabilities to these states can be summed together for these new states. According to this implementation technique, we introduce static vectors that capture the minimum and maximum probabilities of transitions to the target region, with row entries calculated by
$\hat{R}_{\min}(\hat{x},\hat{u},\hat{w})=\sum\limits_{\forall\hat{r}\in\mathcal{T}}\hat{T}_{\min}(\hat{r}\,|\,\hat{x},\hat{u},\hat{w}),\quad\hat{R}_{\max}(\hat{x},\hat{u},\hat{w})=\sum\limits_{\forall\hat{r}\in\mathcal{T}}\hat{T}_{\max}(\hat{r}\,|\,\hat{x},\hat{u},\hat{w}).$
Similarly, for transitions to the avoid region, we define
$\hat{A}_{\min}(\hat{x},\hat{u},\hat{w})=\sum\limits_{\forall\hat{a}\in\mathcal{A}}\hat{T}_{\min}(\hat{a}\,|\,\hat{x},\hat{u},\hat{w}),\quad\hat{A}_{\max}(\hat{x},\hat{u},\hat{w})=\sum\limits_{\forall\hat{a}\in\mathcal{A}}\hat{T}_{\max}(\hat{a}\,|\,\hat{x},\hat{u},\hat{w}).$
Therefore, one can redefine the IMDP depending on the specification as $\hat{\Sigma}=(\hat{X},\hat{U},\hat{W},\hat{T}_{{\min}},\hat{T}_{{\max}},\hat{R}_{{\min}},\hat{R}_{{\max}},\\ \hat{A}_{{\min}},\hat{A}_{{\max}})$. It is worth noting that users do not need to be aware of these additional vectors, as they are automatically computed during the abstraction process from the state labeling, as detailed in Algorithm 1.
3.2. Construction of IMDP
Input: Continious MDP $\Sigma=(X,U,W,T)$ and specification $\psi$
1 Select finite partition of sets $X,U,W$ as $X=\cup^{n_{x}}_{i=0}\textbf{X}^{i}$, $U=\cup^{n_{u}}_{i=0}\textbf{U}^{i}$, $W=\cup^{n_{w}}_{i=0}\textbf{W}^{i}$;
2
3For each $\textbf{X}^{i}$ select a representative point $\hat{x}^{i}\in\textbf{X}^{i}$. Similarly, for each $\textbf{U}^{i}$ and $\textbf{W}^{i}$, select a representative point $\hat{u}^{i}\in\textbf{U}^{i}$ and $\hat{w}^{i}\in\textbf{W}^{i}$;
4 Define $\hat{X}:=\{\hat{x}^{i},i=0,\ldots,n_{x}\}$ as the finite state set of $\hat{\Sigma}$ with input set $\hat{U}:=\{\hat{u}^{i},i=0,\ldots,n_{u}\}$ and disturbance set $\hat{W}:=\{\hat{w}^{i},i=0,\ldots,n_{w}\}$;
5 Define the map $\Xi:X\rightarrow 2^{X}$ that assigns to any $x\in X$, the corresponding partition set it belongs to, i.e., $\Xi(x)=\textbf{X}^{i}$ if $x\in\textbf{X}^{i}$;
6 Label states as safe states $\hat{x}\in\mathcal{S}$, target states $\hat{r}\in\mathcal{T}$, or avoid states $\hat{a}\in\mathcal{A}$, based on desired specification $\psi$;
7 for $\hat{w}_{i}\in\hat{W},i=\{0,\ldots,n_{w}\}$ do
8 for $\hat{u}_{j}\in\hat{U},j=\{0,\ldots,n_{u}\}$ do
9    for $\hat{x}_{k}\in\mathcal{S},k=\{0,\ldots,{n_{s}}\}$  do
10       for $\hat{x}^{\prime}_{l}\in\mathcal{S},l=\{0,\ldots,{n_{s}}\}$  do
11          $\hat{T}_{{\min}}(\hat{x}^{\prime}_{l}\,|\,\hat{x}_{k},\hat{u}_{j},\hat{w}_{i})=\min\limits_{x\in\Xi(\hat{x}_{k})}T(\Xi(\hat{x}^{\prime}_{l})\,|\,x,\hat{u}_{j},\hat{w}_{i})$;
12          $\hat{T}_{{\max}}(\hat{x}^{\prime}_{l}\,|\,\hat{x}_{k},\hat{u}_{j},\hat{w}_{i})=\max\limits_{x\in\Xi(\hat{x}_{k})}T(\Xi(\hat{x}^{\prime}_{l})\,|\,x,\hat{u}_{j},\hat{w}_{i})$;
13         end for
14       $\hat{R}_{{\min}}(\hat{x}_{k},\hat{u}_{j},\hat{w}_{i})=\sum\limits_{\forall\hat{r}\in\mathcal{T}}\min\limits_{x\in\Xi(\hat{x}_{k})}T(\Xi(\hat{r})\,|\,x,\hat{u}_{j},\hat{w}_{i})$;
15       $\hat{R}_{{\max}}(\hat{x}_{k},\hat{u}_{j},\hat{w}_{i})=\sum\limits_{\forall\hat{r}\in\mathcal{T}}\max\limits_{x\in\Xi(\hat{x}_{k})}T(\Xi(\hat{r})\,|\,x,\hat{u}_{j},\hat{w}_{i})$;
16       $\hat{A}_{{\min}}(\hat{x}_{k},\hat{u}_{j},\hat{w}_{i})=\sum\limits_{\forall\hat{a}\in\mathcal{A}}\min\limits_{x\in\Xi(\hat{x}_{k})}T(\Xi(\hat{a})\,|\,x,\hat{u}_{j},\hat{w}_{i})$;
17       $\hat{A}_{{\max}}(\hat{x}_{k},\hat{u}_{j},\hat{w}_{i})=\sum\limits_{\forall\hat{a}\in\mathcal{A}}\max\limits_{x\in\Xi(\hat{x}_{k})}T(\Xi(\hat{a})\,|\,x,\hat{u}_{j},\hat{w}_{i})$;
18
19      end for
20
21   end for
22
23 end for
Output: IMDP $\hat{\Sigma}=(\hat{X},\hat{U},\hat{W},\hat{T}_{{\min}},\hat{T}_{{\max}},\hat{R}_{{\min}},\hat{R}_{{\max}},\hat{A}_{{\min}},\hat{A}_{{\max}})$
Algorithm 1 Serial construction of IMDP
We now describe the approach taken to construct an IMDP $\hat{\Sigma}$, including how to compute minimal and maximal transition probability matrices $\hat{T}_{{\min}}$, $\hat{T}_{{\max}}$, $\hat{R}_{{\min}}$, $\hat{R}_{{\max}}$, $\hat{A}_{{\min}}$, and $\hat{A}_{{\max}}$, which is essential for later-stage verification or controller synthesis procedures. We first present a serial algorithm for the construction of IMDPs, as Algorithm 1, considering stochasticity with normal distributions which are the default noise type in IMPaCT, while providing a custom user-defined distribution to support any arbitrary distributions.
In Algorithm 1, Steps 121-21 - 2 first construct finite partitions and obtain the representative points within $X$, $U$, and $W$. In Step 333, the finite set of states $\hat{X}$, inputs $\hat{U}$, and disturbances $\hat{W}$ are defined, where $n_{x}=|\hat{X}|$, $n_{u}=|\hat{U}|$, and $n_{w}=|\hat{W}|$. In Step 555, we filter the states based on their labeling for the specification $\psi$. In Steps 101110-1110 - 11, the transition probability matrices $\hat{T}_{{\min}}$ and $\hat{T}_{{\max}}$ are computed by iterating through all the combinations of safe states $\hat{x}\in\mathcal{S}$, where $n_{s}=|\mathcal{S}|$, inputs $\hat{u}\in\hat{U}$, and disturbances $\hat{w}\in\hat{W}$. To do so, a nonlinear optimization is required to find $x\in\Xi(\hat{x})$ corresponding to either the lower bound or upper bound probability of transitioning from $\Xi(\hat{x})$ to $\Xi(\hat{x}^{\prime})$. Similarly, Steps 131613-1613 - 16 perform the same procedure for one-step transitions going to either the target or avoid regions. The output of the algorithm is therefore IMDP $\hat{\Sigma}=(\hat{X},\hat{U},\hat{W},\hat{T}_{{\min}},\hat{T}_{{\max}},\hat{R}_{{\min}},\hat{R}_{{\max}},\hat{A}_{{\min}},\hat{A}_{{\max}})$.
Remark 3.3.
One can readily construct interval Markov chains (IMCs) for verification purposes via Algorithm 1 by setting the control input zero within the concrete dynamics, i.e., $\Sigma=(X,W,T)$.
3.3. Complexity Analysis for Serial Construction of IMDP
To conduct a computational complexity analysis for the serial construction of IMDP, we adopt a slight abuse of notation by defining $n_{x_{i}}$, $n_{u_{j}}$, and $n_{w_{k}}$ as the dimension-wise counts of partitions within $\hat{X}$, $\hat{U}$, and $\hat{W}$, respectively, where $i=1,\ldots,n$, $j=1,\ldots,m$, and $k=1,\ldots,p$. The computational complexity of IMDP construction comprises two main components. The first, often termed the curse of dimensionality, describes the exponential rise in computational time concerning system dimensions, expressed as $\mathcal{O}(2d)$, where $d=(n_{x_{i}}^{n}\times n_{u_{j}}^{m}\times n_{w_{k}}^{p})\times n_{x_{i}}^{n}$. The value $d$ represents the total count of rows in the static transition matrix, i.e., $n_{x}\times n_{x}\times n_{u}$, multiplied by the number of columns, i.e., $n_{x}$. Given that the construction of IMDP abstraction involves both lower bound and upper bound transition matrices $\hat{T}_{\min}$ and $\hat{T}_{\max}$, respectively, the computational complexity $d$ is doubled.
The second part pertains to the computational complexity of the nonlinear optimization within Steps 101110-1110 - 11 in Algorithm 1, represented as $\mathcal{O}({\kappa})$. The level of complexity $\kappa$ hinges on the algorithm chosen by the user, detailed in Section 4.3. Consequently, the overall complexity of Algorithm 1 amounts to $\mathcal{O}({2\kappa}d)$. This accounts for both the space explosion ($d$) due to the system dimensions and the computational load ($\kappa$) associated with nonlinear optimization.
4. Parallel Construction of IMDP
In this section, we describe how we update Algorithm 1 for parallel processing as offered by IMPaCT. We also here present the user functions of our tool corresponding to various subsections. Firstly, an IMDP object should be constructed by defining the dimensions of state, input and disturbance:
1IMDP(const int x, const int u, const int w);
Listing 1: Creating IMDP object.
4.1. Setting Noise Distribution
IMPaCT by default supports normal distributions but also allows a user-defined distribution to be provided:
1enum class NoiseType {NORMAL, CUSTOM};
Listing 2: Define noise distribution.
We consider here an i.i.d. noise where the tool receives a vector of standard deviations. In this setup, the covariance matrix is diagonal, and each diagonal entry represents the variance:
1void setStdDev(vec sig);
2void setNoise(NoiseType n, bool diagonal = true);
Listing 3: Diagonal covariance matrix.
IMPaCT also accommodates full covariance matrix, with nonzero off-diagonal elements, when provided with the inverse covariance matrix and its determinant. For multi-dimensional noise distributions, Monte Carlo integration is employed instead of a closed-form solution, offering advantages for higher dimension integrals [NB99], requiring the user to specify the number of samples for the integration process. We utilize the GNU Scientific Library for this integration process [Gou09]. In a broader scope, matrix and vector manipulations are implemented using the C++ library Armadillo [SC16, SC18]:
1void setInvCovDet(mat inv_cov, double det);
2void setNoise(NoiseType n, bool diagonal, size_t monte_carlo_samples);
Listing 4: Full covariance matrix.
We facilitate the utilization of a user-defined custom distribution, prompting the user to define the specific desired distribution. The code will provide a struct customParams containing data for the variables state_start $\hat{x}$, input $\hat{u}$, disturb $\hat{w}$, eta $\eta_{x}$, mean $f(\hat{x},\hat{u},\hat{w})$, lb lower bound of the integration, ub upper bound of the integration, and dynamics{1,2,3} dynamics with numbers representing the number of function parameters. The customPDF function intended for integration can be coded within the file src/custom.cpp, and the parameters for the Monte Carlo integration–CUSTOM noise and the number of samples–can be configured using:
1void setNoise(NoiseType n);
2void setCustomDistribution(size_t monte_carlo_samples);
Listing 5: Custom noise distributions.
4.2. Parallel Construction of Transition Probability Matrices
We propose in Algorithm 2 the parallel approach for constructing the transition matrices $\hat{T}_{{\min}}$, $\hat{T}_{{\max}}$, $\hat{R}_{{\min}}$, $\hat{R}_{{\max}}$, $\hat{A}_{{\min}}$, and $\hat{A}_{{\max}}$. Steps 151-51 - 5 are equivalent to Algorithm 1. In IMPaCT, these Steps 151-51 - 5, can be computed using the following user functions. The dimensions and representative points within each space can be specified by setting the lower bounds, upper bounds, and discretization parameters. As elaborated in Subsection 3.1, we consider states within the state space as safe states by default, constituting the safe region. Subsequently, the target and avoid regions can be delineated by employing Boolean expressions that assess the state space. These expressions facilitate the relabeling of states, transitioning them from safe states to either target or avoid states based on desired specifications:
1void setStateSpace(vec lb, vec ub, vec eta);
2void setInputSpace(vec lb, vec ub, vec eta);
3void setDisturbSpace(vec lb, vec ub, vec eta);
4void setTargetSpace(const function<bool(const vec&)>& separate_condition, bool remove);
5void setAvoidSpace(const function<bool(const vec&)>& separate_condition, bool remove);
6void setTargetAvoidSpace(const function<bool(const vec&)>& target_condition,const function<bool(const vec&)>& avoid_condition, bool remove);
Listing 6: Construction of $\hat{X},\hat{U},\hat{W},\mathcal{T}$, and $\mathcal{A}$.
Input: MDP $\Sigma=(X,U,W,T)$ and specification $\psi$
1 Select finite partition of sets $X,U,W$ as: $X=\cup^{n_{x}}_{i=0}\textbf{X}^{i}$, $U=\cup^{n_{u}}_{i=0}\textbf{U}^{i}$, $W=\cup^{n_{w}}_{i=0}\textbf{W}^{i}$;
2
3For each partition $\textbf{X}^{i}$ select a representative point $\hat{x}^{i}\in\textbf{X}^{i}$ used to describe the partition, e.g., centre of the partition. Similarly for $\textbf{U}^{i}$ and $\textbf{W}^{i}$, select representative points $\hat{u}^{i}\in\textbf{U}^{i}$ and $\hat{w}^{i}\in\textbf{W}^{i}$;
4 Define $\hat{X}:=\{\hat{x}^{i},i=0,\ldots,n_{x}\}$ as the finite state set of $\hat{\Sigma}$ with input set $\hat{U}:=\{\hat{u}^{i},i=0,\ldots,n_{u}\}$ and disturbance set $\hat{W}:=\{\hat{w}^{i},i=0,\ldots,n_{w}\}$;
5 Define the map $\Xi:X\rightarrow 2^{X}$ that assigns to any $x\in X$, the corresponding partition set it belongs to, i.e., $\Xi(x)=\textbf{X}^{i}$ if $x\in\textbf{X}^{i}$;
6 Relabel states as safe states $\hat{x}\in\mathcal{S}$, target states $\hat{r}\in\mathcal{T}$, or avoid states $\hat{a}\in\mathcal{A}$ based on desired specification $\psi$;
7 forall combinations of $\hat{x}_{k}\in\mathcal{S},\hat{u}_{j}\in\hat{U},$ and $\hat{w}_{i}\in\hat{W}$ in parallel do
8 for $\hat{x}^{\prime}_{l}\in\mathcal{S},l=\{0,\ldots,n_{x}\}$ do
9    $\hat{T}_{{\min}}(\hat{x}^{\prime}_{l}|\hat{x}_{k},\hat{u}_{j},\hat{w}_{i})=\min\limits_{x\in\Xi(\hat{x}_{k})}T(\Xi(\hat{x}^{\prime}_{l})|x,\hat{u}_{j},\hat{w}_{i})$;
10    $\hat{T}_{{\max}}(\hat{x}^{\prime}_{l}|\hat{x}_{k},\hat{u}_{j},\hat{w}_{i})=\max\limits_{x\in\Xi(\hat{x}_{k})}T(\Xi(\hat{x}^{\prime}_{l})|x,\hat{u}_{j},\hat{w}_{i})$;
11   end for
12 $\hat{R}_{{\min}}(\hat{x}_{k},\hat{u}_{j},\hat{w}_{i})=\sum\limits_{\forall\hat{r}\in\mathcal{T}}\min\limits_{x\in\Xi(\hat{x}_{k})}T(\Xi(\hat{r})|x,\hat{u}_{j},\hat{w}_{i})$;
13 $\hat{R}_{{\max}}(\hat{x}_{k},\hat{u}_{j},\hat{w}_{i})=\sum\limits_{\forall\hat{r}\in\mathcal{T}}\max\limits_{x\in\Xi(\hat{x}_{k})}T(\Xi(\hat{r})|x,\hat{u}_{j},\hat{w}_{i})$;
14 $\hat{A}_{{\min}}(\hat{x}_{k},\hat{u}_{j},\hat{w}_{i})=\sum\limits_{\forall\hat{a}\in\mathcal{A}}\min\limits_{x\in\Xi(\hat{x}_{k})}T(\Xi(\hat{a})|x,\hat{u}_{j},\hat{w}_{i})$;
15 $\hat{A}_{{\max}}(\hat{x}_{k},\hat{u}_{j},\hat{w}_{i})=\sum\limits_{\forall\hat{a}\in\mathcal{A}}\max\limits_{x\in\Xi(\hat{x}_{k})}T(\Xi(\hat{a})|x,\hat{u}_{j},\hat{w}_{i})$;
16
17 end forall
Output: IMDP $\hat{\Sigma}=(\hat{X},\hat{U},\hat{W},\hat{T}_{{\min}},\hat{T}_{{\max}},\hat{R}_{{\min}},\hat{R}_{{\max}},\hat{A}_{{\min}},\hat{A}_{{\max}})$
Algorithm 2 Parallel Construction of IMDP
In Steps 898-98 - 9, transition probabilities are computed in parallel for the value $x\in\Xi(\hat{x})$ corresponding to either the minimal (lower bound) or maximal (upper bound) probability of transitioning from $\Xi(\hat{x})$ to $\Xi(\hat{x}^{\prime})$. In Step 111211-1211 - 12, target vector bounds $\hat{R}_{{\min}}$ and $\hat{R}_{{\max}}$ calculate the one-step transition probability of the state $\hat{x}$ transitioning to $\mathcal{T}$. Similarly, avoid vector bounds $\hat{A}_{{\min}}$ and $\hat{A}_{{\max}}$ calculate the transition probability of the state $\hat{x}$ transitioning to $\mathcal{A}$. For each state-input-disturbance triple, we sum the probabilities together resulting in the respective target vector or avoid vector. In IMPaCT, Steps 898-98 - 9 and 111411-1411 - 14, can be computed using the following function calls:
1void minTransitionMatrix();
2void maxTransitionMatrix();
3void minTargetTransitionVector();
4void maxTargetTransitionVector();
5void minAvoidTransitionVector();
6void maxAvoidTransitionVector();
Listing 7: Abstraction of transition matrices.
Remark 4.1.
In IMPaCT, given that the state space is bounded, some case studies may involve potential transitions outside the defined state space. This is captured in the functions minAvoidTransitionVector() as $\hat{A}_{\min}(\hat{x},\hat{u},\hat{w})=1-\hat{T}_{\max}(\hat{X}|\hat{x},\hat{u},\hat{w})$ and with a similar construction for maxAvoidTransitionVector(). These functions also consider the minimal and maximal transition probabilities to any additional avoid states that are inside the state space by summing them to $\hat{A}_{\min}(\hat{x},\hat{u},\hat{w})$ and $\hat{A}_{\max}(\hat{x},\hat{u},\hat{w})$, respectively.
4.3. Complexity Analysis for Parallel Construction of IMDP
Running these steps in parallel enhances the performance of solving the abstraction and synthesis problems. Using $\mathcal{O}(\kappa)$ to represent the optimization complexity, the complexity of Algorithm 2 is therefore $\mathcal{O}(\frac{2\kappa d}{\textsc{threads}}),$ where $d=(n_{x_{i}}^{n}\times n_{u_{j}}^{m}\times n_{w_{k}}^{p})\times n_{x_{i}}^{n}$, and threads is the number of parallel threads running. We implement the optimization algorithms using NLopt [Joh07], by default we select nlopt::LN_SBPLX [Row90], which is a variant of the derivative-free Nelder-Mead Simplex algorithm. It is worth highlighting that the chosen algorithm in IMPaCT can be readily swapped to any other nonlinear optimization algorithm from NLopt [^3] https://nlopt.readthedocs.io/en/latest/NLopt_\__Algorithms/ using the following function:
1void setAlgorithm(nlopt::algorithm alg);
Listing 8: Set NLopt algorithm.
The primary bottleneck in the IMDP abstraction arises from the necessity of computing two transition probability matrices, coupled with the additional computational load of $\min$/$\max$ optimization. To enhance performance, by computing the matrix $\hat{T}_{{\max}}$ first, the computations required for $\hat{T}_{{\min}}$ can be reduced. This is due to the fact that for any matrix entry $\hat{T}_{{\max}}(i,j)=0$, one has $\hat{T}_{{\min}}(i,j)=0$. The sparser the matrix, the more efficient the abstraction computation; we refer to this as the low-cost abstraction. This speed up can be used by calling the following functions for the transition matrices and the target transition vectors, respectively:
1void transitionMatrixBounds();
2void targetTransitionMatrixBounds();
Listing 9: Low-cost abstractions.
5. Controller Synthesis with Interval Iteration
We now synthesize a controller, via constructed IMDP, to enforce infinite-horizon properties over dt-SCS. When dealing with infinite-horizon problems, the interval iteration algorithm is capable of providing convergence guarantees unlike the common value iteration [HM18]. In particular, the interval iteration algorithm converges to an under-approximation and over-approximation of the satisfaction probability associated with a temporal logic specification. The trade-off for achieving such a guarantee is the doubled computational load compared to the value iteration algorithm.
In essence, the interval iteration algorithm iterates over two Bellman equations simultaneously; one assuming an initial probability vector of zeros, denoted by $V_{0}=\textbf{0}_{n}$, and the other an initial vector of ones, represented as $V_{1}=\textbf{1}_{n}$. When calculating $V_{0}^{\prime}$ and $V_{1}^{\prime}$ as the next iteration step, a dynamic program with decision variables $\hat{R}$, $\hat{A}$, and $\hat{T}$ needs to be solved [HM18]. In finite MDP construction, a state within a partition is predetermined beforehand, and the probability transitions to all other partitions are computed based on this state. However, in IMDP construction, each partition might utilize a distinct state when calculating probability transitions to another partition, contingent upon the destination partition, i.e., the partition where the system lands after a one-step evolution. Consequently, solving the dynamic program can be seen as akin to determining a partition’s state that could universally serve for computing probability transitions to other partitions. This solution is referred to as a feasible distribution. In particular, the feasible distribution is determined to optimize (either minimize or maximize) the decision variables given the desired property of interest. Subsequently, the corresponding optimization is minimized with respect to the disturbance $\hat{w}$ and maximized concerning the input $\hat{u}$:
$\max\limits_{\hat{u}\in\hat{U}}\min\limits_{\hat{w}\in\hat{W}}\underset{\hat{R},\hat{A},\hat{T}}{\text{optimize}}~\delta_{1}\hat{R}(\hat{x},\hat{u},\hat{w})+\delta_{2}\hat{A}(\hat{x},\hat{u},\hat{w})+\sum\limits_{\forall\hat{x}^{\prime}\in\hat{X}}\delta_{3}(\hat{x}^{\prime})\hat{T}(\hat{x}^{\prime}|\hat{x},\hat{u},\hat{w})$ (5.1)
with weight functions $\delta_{1}$, $\delta_{2}$, and $\delta_{3}(\cdot)$, and being subject to the following constraints [HM18]:
$\hat{T}_{\min}(\hat{x}^{\prime}|\hat{x},\hat{u},\hat{w})\leq\hat{T}(\hat{x}^{\prime}|\hat{x},\hat{u},\hat{w})\leq\hat{T}_{\max}(\hat{x}^{\prime}|\hat{x},\hat{u},\hat{w}),$
$\hat{R}_{\min}(\hat{x},\hat{u},\hat{w})\leq\hat{R}(\hat{x},\hat{u},\hat{w})\leq\hat{R}_{\max}(\hat{x},\hat{u},\hat{w}),$
$\hat{A}_{\min}(\hat{x},\hat{u},\hat{w})\leq\hat{A}(\hat{x},\hat{u},\hat{w})\leq\hat{A}_{\max}(\hat{x},\hat{u},\hat{w}),$
$\hat{R}(\hat{x},\hat{u},\hat{w})+\hat{A}(\hat{x},\hat{u},\hat{w})+\sum\limits_{\forall\hat{x}^{\prime}\in\hat{X}}\hat{T}(\hat{x}^{\prime}|\hat{x},\hat{u},\hat{w})=1.$
In particular, $\hat{R}$ mimics $s^{+}$ and $\hat{A}$ mimics $s^{-}$ in the formulation of [HM18]. The corresponding weights are attributed to the dynamic program and are updated at each iteration step. The controller synthesis here uses a linear program (LP) to converge to an optimal solution, where the weights are the probability of satisfying the specification and $\delta_{3}$ is derived from either $V_{0}$ or $V_{1}$. Specifically, the optimization in (5.1) takes into account the probability from transitioning from a partition with representative state $\hat{x}$ to another partition with representative state $\hat{x}^{\prime}$ multiplied by some weight. This weight corresponds to the probability of satisfying the specification when initiating from a state within the partition indicated by $\hat{x}^{\prime}$. The affine part of the optimization depends on the specification for the weights attributed to $\hat{R}$ and $\hat{A}$. For reachability specifications, $\delta_{1}=1$ and $\delta_{2}=0$, whereas for safety, the inverse holds true. Additionally, the probability of fulfilling the specification complements the resulting optimal values for $V_{0}$ and $V_{1}$.
When the dynamic program is solved and the optimal feasible solutions $\hat{T}$, $\hat{R}$, and $\hat{A}$ are found, the interval iteration algorithm solves the two equations
$\begin{cases}V_{0}^{\prime}=\delta_{1}\hat{R}+\delta_{2}\hat{A}+\hat{T}V_{0},\\ V_{1}^{\prime}=\delta_{1}\hat{R}+\delta_{2}\hat{A}+\hat{T}V_{1},\end{cases}$ (5.2)
to find the new probabilities of satisfying the specification for the state-input-disturbance triples. Prior to the subsequent iteration, $V_{0}$ and $V_{1}$ are, respectively, updated to $V_{0}^{\prime}$ and $V_{1}^{\prime}$. The interval iteration algorithm terminates when the two vectors converge, $\lVert V_{1}-V_{0}\rVert_{\infty}\leq\varepsilon$, where $\varepsilon$ is set by default to a small enough threshold.
The resulting controller $\mathcal{C}=(\hat{X},\pi,\mathds{P}_{\psi_{\min}},\mathds{P}_{\psi_{\max}})$ consists of a lookup table with the optimal control policy $\pi$ for each representative point $\hat{x}\in\mathcal{S}$, the minimal probability of satisfying the specification $\mathds{P}_{\psi_{\min}}$ to and the maximal probability of satisfying the specification $\mathds{P}_{\psi_{\max}}$. When synthesizing the controller, either $\mathds{P}_{\psi_{\min}}$ or $\mathds{P}_{\psi_{\max}}$ is prioritized for finding $\pi$, giving a pessimistic or optimistic control policy, respectively. In particular, a pessimistic policy is when the policy $\pi$ is optimized using (5.1) that minimizes $\hat{A},\hat{R}$ and $\hat{T}$, while an optimistic policy maximizes $\hat{A},\hat{R}$ and $\hat{T}$ [DLML23]. The policy $\pi$ is then fixed when calculating the other bound, e.g., $\pi$ synthesized from calculations on $\mathds{P}_{\psi_{\min}}$ is fixed when finding $\mathds{P}_{\psi_{\max}}$.
The underlying synthesis technique is equivalent to a three-and-a-half player game with a max-min optimization problem, where the input and disturbance are two players. In particular, treating the disturbance as an adversary to the system entails minimizing the probability concerning the disturbance while maximizing it concerning the control inputs. The third player represents the range across the targeted partition addressed by the optimization program, acting as an adversary in computing $\mathds{P}_{\psi_{\min}}$ and an ally when computing $\mathds{P}_{\psi_{\max}}$.
5.1. Safety Specifications
Safety specifications seem to be typically the simplest of the three specification types handled by IMPaCT. However, the interval iteration algorithm for safety specifications is not entirely straightforward. In particular, for controller synthesis via finite MDPs, value iteration is used where the transition probability matrix $\hat{T}$ is multiplied by a vector of ones $V_{1}$ at each iteration backward over time, e.g. $V_{1}^{\prime}=\hat{T}V_{1}$. In contrast, interval iteration in IMDP requires a static affine vector (i.e., either $\hat{A}$ or $\hat{R}$), otherwise $V_{0}^{\prime}$ will always remain $\textbf{0}_{n}$ as $V_{0}^{\prime}=\hat{T}V_{0}$ with $V_{0}=\textbf{0}_{n}$, and $V_{1}^{\prime}=\hat{T}V_{1}$ with $V_{0}=\textbf{1}_{n}$ at the start of the Bellman equation. Hence, the algorithm cannot converge, except when $V_{1}$ converges to $\textbf{0}_{n}$, indicating zero safety guarantee.
To resolve this challenge, given that safety and reachability are complement of each other, we fulfill our safety specification by solving a reachability problem over the complement of the safe set. Logically speaking, $\square\mathcal{S}$ is equivalent to $\neg\lozenge\neg\mathcal{S}$, where $\neg$ refers to the LTL term “not”. Therefore, the dynamic program to be solved is (5.1), where $\delta_{1}=0$, $\delta_{2}=1$, and $\delta_{3}=V_{1}$, and the set $\mathcal{T}=\emptyset$. Additionally, we minimize with respect to the input $\hat{u}$ and maximize with respect to the disturbance $\hat{w}$. This approach stems from the intuition behind solving a reachability problem toward an undesired area. Consequently, we aim to select the input least likely to reach the avoid region, which equates to choosing the input most likely to remain within the safe region. Similarly, the adversary selects the disturbance most likely to reach the avoid region. When the algorithm converges, the complement is taken to give the resulting $\mathds{P}_{\psi_{\min}}$ and $\mathds{P}_{\psi_{\max}}$. It is notable that for safety specifications over an infinite time horizon, it is not uncommon for the solution to converge to the trivial result where $\mathds{P}_{\psi_{\min}}=0$ and $\mathds{P}_{\psi_{\max}}=1$, or for both bounds to converge to 0.
Input: IMDP $\hat{\Sigma}$, stopping condition $\varepsilon$
1 $V_{0}:=\textbf{0}_{n}$, $V_{1}:=\textbf{1}_{n}$, $\pi:=\textbf{0}_{n}$;
2 while $\lVert V_{1}-V_{0}\rVert{{}_{\infty}}>\varepsilon$ do
3 for $\hat{x}_{k}\in\mathcal{S},k=\{0,\ldots,|\mathcal{S}|\}$ do
4    Solve max LP (5.1) for $\hat{A}$ and $\hat{T}$;
5    $\pi(\hat{x}_{k}):=\operatorname*{argmin}\limits_{\hat{u}\in\hat{U}}\max\limits_{\hat{w}\in\hat{W}}{\max\limits_{\hat{A},\hat{T}}}\{\hat{A}(\hat{x}_{k},\cdot,\cdot)+\sum\limits_{\forall\hat{x}^{\prime}\in\mathcal{S}}\hat{T}(\cdot|\hat{x}_{k},\cdot,\cdot)V_{0}(\hat{x}^{\prime})\}$;
6    $V_{0}^{\prime}(\hat{x}_{k}):=\min\limits_{\hat{u}\in\hat{U}}\max\limits_{\hat{w}\in\hat{W}}{\max\limits_{\hat{A},\hat{T}}}\{\hat{A}(\hat{x}_{k},\cdot,\cdot)+\sum\limits_{\forall\hat{x}^{\prime}\in\mathcal{S}}\hat{T}(\cdot|\hat{x}_{k},\cdot,\cdot)V_{0}(\hat{x}^{\prime})\}$;
7    $V_{1}^{\prime}(\hat{x}_{k}):=\min\limits_{\hat{u}\in\hat{U}}\max\limits_{\hat{w}\in\hat{W}}{\max\limits_{\hat{A},\hat{T}}}\{\hat{A}(\hat{x}_{k},\cdot,\cdot)+\sum\limits_{\forall\hat{x}^{\prime}\in\mathcal{S}}\hat{T}(\cdot|\hat{x}_{k},\cdot,\cdot)V_{1}(\hat{x}^{\prime})\}$;
8
9   end for
10 $V_{0}:=V_{0}^{\prime},~V_{1}:=V_{1}^{\prime}$;
11
12 end while
[…]
14 $V_{0}:=\textbf{0}_{n}$, $V_{1}:=\textbf{1}_{n}$;
15
16while $\lVert V_{1}-V_{0}\rVert{{}_{\infty}}>\varepsilon$ do
17 for $\hat{x}_{k}\in\mathcal{S},k=\{0,\ldots,|\mathcal{S}|\}$ do
18    Solve min LP (5.1) for $\hat{A}$ and $\hat{T}$;
19      fix $\hat{u}$ from $\pi(\hat{x}_{k})$;
20    $V_{0}^{\prime}(\hat{x}_{k}):=\max\limits_{\hat{w}\in\hat{W}}{\min\limits_{\hat{A},\hat{T}}}\{\hat{A}(\hat{x}_{k},\hat{u},\cdot)+\sum\limits_{\forall\hat{x}^{\prime}\in\mathcal{S}}\hat{T}(\cdot|\hat{x}_{k},\hat{u},\cdot)V_{0}(\hat{x}^{\prime})\}$;
21    $V_{1}^{\prime}(\hat{x}_{k}):=\max\limits_{\hat{w}\in\hat{W}}{\min\limits_{\hat{A},\hat{T}}}\{\hat{A}(\hat{x}_{k},\hat{u},\cdot)+\sum\limits_{\forall\hat{x}^{\prime}\in\mathcal{S}}\hat{T}(\cdot|\hat{x}_{k},\hat{u},\cdot)V_{1}(\hat{x}^{\prime})\}$;
22
23   end for
24 $V_{0}:=V_{0}^{\prime},~V_{1}:=V_{1}^{\prime}$;
25
26 end while
[…]
Output: controller $\mathcal{C}=(\mathcal{S},\pi,\mathds{P}_{\psi_{\min}},\mathds{P}_{\psi_{\max}})$
Algorithm 3 Serial (Pessimistic) Safety Controller Synthesis
In Algorithm 3, controller synthesis for a pessimistic safety specification in serial is displayed. Steps 676-76 - 7, and Steps 171817-1817 - 18 show the interval iteration algorithm where the first loops find the optimal pessimistic control policy $\pi$ using the lower bound and the second loops fix the control policy to find the IMDP upper bound. Notice that when calculating $\mathds{P}_{\psi_{\min}}$, we solve the max LP, and for $\mathds{P}_{\psi_{\max}}$ we solve the min LP. Step 121212 and Step 222222 show the complement is taken, giving us the safety probability bounds.
Input: IMDP $\hat{\Sigma}$, stopping condition $\varepsilon$
1 $V_{0}:=\textbf{0}_{n}$, $V_{1}:=\textbf{1}_{n}$, $\pi:=\textbf{0}_{n}$;
2 while $\lVert V_{1}-V_{0}\rVert{{}_{\infty}}>\varepsilon$ do
3 for $\hat{x}_{k}\in\mathcal{S},k=\{0,\ldots,|\mathcal{S}|\}$ do
4    Solve min LP (5.1) for $\hat{R}$ and $\hat{T}$;
5    $\pi(\hat{x}_{k}):=\operatorname*{argmax}\limits_{\hat{u}\in\hat{U}}\min\limits_{\hat{w}\in\hat{W}}{\min\limits_{\hat{R},\hat{T}}}\{\hat{R}(\hat{x}_{k},\cdot,\cdot)+\sum\limits_{\forall\hat{x}^{\prime}\in\hat{X}}\hat{T}(\cdot|\hat{x}_{k},\cdot,\cdot)V_{0}(\hat{x}^{\prime})\}$;
6    $V_{0}^{\prime}(\hat{x}_{k}):=\max\limits_{\hat{u}\in\hat{U}}\min\limits_{\hat{w}\in\hat{W}}{\min\limits_{\hat{R},\hat{T}}}\{\hat{R}(\hat{x}_{k},\cdot,\cdot)+\sum\limits_{\forall\hat{x}^{\prime}\in\hat{X}}\hat{T}(\cdot|\hat{x}_{k},\cdot,\cdot)V_{0}(\hat{x}^{\prime})\}$;
7    $V_{1}^{\prime}(\hat{x}_{k}):=\max\limits_{\hat{u}\in\hat{U}}\min\limits_{\hat{w}\in\hat{W}}{\min\limits_{\hat{R},\hat{T}}}\{\hat{R}(\hat{x}_{k},\cdot,\cdot)+\sum\limits_{\forall\hat{x}^{\prime}\in\hat{X}}\hat{T}(\cdot|\hat{x}_{k},\cdot,\cdot)V_{1}(\hat{x}^{\prime})\}$;
8
9   end for
10 $V_{0}:=V_{0}^{\prime},~V_{1}:=V_{1}^{\prime}$;
11
12 end while
[…]
14 $V_{0}:=\textbf{0}_{n}$, $V_{1}:=\textbf{1}_{n}$;
15 while $\lVert V_{1}-V_{0}\rVert{{}_{\infty}}>\varepsilon$ do
16 for $\hat{x}_{k}\in\mathcal{S},k=\{0,\ldots,|\mathcal{S}|\}$ do
17    Solve max LP (5.1) for $\hat{R}$ and $\hat{T}$;
18      fix $\hat{u}$ from $\pi(\hat{x}_{k})$;
19    $V_{0}^{\prime}(\hat{x}_{k}):=\min\limits_{\hat{w}\in\hat{W}}{\max\limits_{\hat{R},\hat{T}}}\{\hat{R}(\hat{x}_{k},\hat{u},\cdot)+\sum\limits_{\forall\hat{x}^{\prime}\in\hat{X}}\hat{T}(\cdot|\hat{x}_{k},\hat{u},\cdot)V_{0}(\hat{x}^{\prime})\}$;
20    $V_{1}^{\prime}(\hat{x}_{k}):=\min\limits_{\hat{w}\in\hat{W}}{\max\limits_{\hat{R},\hat{T}}}\{\hat{R}(\hat{x}_{k},\hat{u},\cdot)+\sum\limits_{\forall\hat{x}^{\prime}\in\hat{X}}\hat{T}(\cdot|\hat{x}_{k},\hat{u},\cdot)V_{1}(\hat{x}^{\prime})\}$;
21
22   end for
23 $V_{0}:=V_{0}^{\prime},~V_{1}:=V_{1}^{\prime}$;
24
25 end while
[…]
Output: controller $\mathcal{C}=(\mathcal{S},\pi,\mathds{P}_{\psi_{\min}},\mathds{P}_{\psi_{\max}})$
Algorithm 4 Serial (Pessimistic) Reach-Avoid Controller Synthesis
5.2. Reachability and Reach-while-Avoid Specifications
Reachability specifications typically introduce additional complexities compared to safety specifications, primarily due to the set of target states that the control policy must guide the system toward. Specifically, states that are not in the target (or avoid) region are considered safe, but the system should make progress towards the target set $\mathcal{T}$. Interval iteration for reachability is relatively straightforward. In our setting, $\hat{R}$ represents the transition probability of a state to the target region $\mathcal{T}$.
Reach-while-avoid specifications consider the case where $\hat{a}\in\mathcal{A}$ in the state space should be avoided. Then the dynamic program to be solved is (5.1), where $\delta_{1}=1$, $\delta_{2}=0$, and $\delta_{3}=V_{1}$. When the algorithm converges, the resulting solution is $\mathds{P}_{\psi_{\min}}$ and $\mathds{P}_{\psi_{\max}}$ for each bound. It is worth noting that the synthesis of a reachability specification is equivalent to a reach-while-avoid specification where the avoid region $\mathcal{A}=\emptyset$.
A serial pessimistic reach-while-avoid controller approach is outlined in Algorithm 4, in which the interval iteration occurs in Steps 676-76 - 7 and Steps 171817-1817 - 18. The first while loop iterates to synthesize a control policy $\pi$, and the second while loop uses $\pi$ by fixing the input $\hat{u}$ for each state $\hat{x}_{k}$. Steps 333 and 141414 show the reach-avoid specification where the avoid set is removed from the state space leaving only $\hat{x}\in\mathcal{S}$. As earlier mentioned, the algorithm remains equivalent for the simpler reachability specification when $\mathcal{A}=\emptyset$. Steps 111111 and 222222 store the minimal and maximal probabilities of satisfaction corresponding to the policy $\pi$. Compared with the safety algorithm, no complement is required when calculating those probabilities.
Remark 5.1.
IMPaCT can also support finite-horizon specifications within IMDP construction by replacing the while loop in Algorithms 3 and 4 with a for loop over a finite number of time intervals. If this scenario applies, we still calculate each bound separately, fixing the policy $\pi$ for the second bound. In this case, interval iteration can be reduced to value iteration, enabling the solution of finite-horizon problems, albeit with the trade-off of sacrificing convergence guarantees [HM18].
Remark 5.2.
In certain scenarios, such as the presence of absorbing states beyond those already outlined in the target and/or avoid regions, the two bounds of the interval iteration algorithm might not converge. In this case, calculating the number of time intervals $k$ for which each independent bound of the interval iteration algorithm converges to itself (with a small enough threshold) can be beneficial. The larger $k$ value can then be employed by the value iteration algorithm as its time horizon to hopefully ascertain a converged solution, ensuring the correctness of convergence. IMPaCT offers detailed guidance on the required steps to follow in such cases, including the example ex_load_safe.
6. Parallel Controller Synthesis with Interval Iteration
As offered in our approach, we can improve the computational efficiency of solving synthesis problems to acquire the resulting controller $\mathcal{C}=(\mathcal{S},\pi,\mathds{P}_{\psi_{\min}},\mathds{P}_{\psi_{\max}})$ by using parallel processing. The synthesis includes solving several dynamic programs and two interval iteration algorithms. We use the GNU linear programming kit [Mak08] for solving the linear programs.
Input: IMDP $\hat{\Sigma}$, stopping condition $\varepsilon$
1 $V_{0}:=\textbf{0}_{n}$, $V_{1}:=\textbf{1}_{n}$, $\pi:=\textbf{0}_{n}$;
2 while $\lVert V_{1}-V_{0}\rVert{{}_{\infty}}>\varepsilon$ do
3 for $x\in\mathcal{S}$ in parallel do
4    Solve max LP (5.1) for $\hat{A}$ and $\hat{T}$;
5    $\pi(\hat{x}_{k}):=\operatorname*{argmin}\limits_{\hat{u}\in\hat{U}}\max\limits_{\hat{w}\in\hat{W}}{\max\limits_{\hat{A},\hat{T}}}\{\hat{A}(\hat{x}_{k},\cdot,\cdot)+\sum\limits_{\forall\hat{x}^{\prime}\in\hat{X}}\hat{T}(\cdot|\hat{x}_{k},\cdot,\cdot)V_{0}(\hat{x}^{\prime})\}$;
6    $V_{0}^{\prime}(\hat{x}_{k}):=\min\limits_{\hat{u}\in\hat{U}}\max\limits_{\hat{w}\in\hat{W}}{\max\limits_{\hat{A},\hat{T}}}\{\hat{A}(\hat{x}_{k},\cdot,\cdot)+\sum\limits_{\forall\hat{x}^{\prime}\in\hat{X}}\hat{T}(\cdot|\hat{x}_{k},\cdot,\cdot)V_{0}(\hat{x}^{\prime})\}$;
7    $V_{1}^{\prime}(\hat{x}_{k}):=\min\limits_{\hat{u}\in\hat{U}}\max\limits_{\hat{w}\in\hat{W}}{\max\limits_{\hat{A},\hat{T}}}\{\hat{A}(\hat{x}_{k},\cdot,\cdot)+\sum\limits_{\forall\hat{x}^{\prime}\in\hat{X}}\hat{T}(\cdot|\hat{x}_{k},\cdot,\cdot)V_{1}(\hat{x}^{\prime})\}$;
8
9   end for
10 $V_{0}:=V_{0}^{\prime},~V_{1}:=V_{1}^{\prime}$;
11
12 end while
[…]
14 $V_{0}:=\textbf{0}_{n}$, $V_{1}:=\textbf{1}_{n}$;
15
16while $\lVert V_{1}-V_{0}\rVert{{}_{\infty}}>\varepsilon$ do
17 for $x\in\mathcal{S}$ in parallel do
18    Solve min LP (5.1) for $\hat{A}$ and $\hat{T}$;
19      fix $\hat{u}$ from $\pi(\hat{x}_{k})$;
20    $V_{0}^{\prime}(\hat{x}_{k}):=\max\limits_{\hat{w}\in\hat{W}}{\min\limits_{\hat{A},\hat{T}}}\{\hat{A}(\hat{x}_{k},\hat{u},\cdot)+\sum\limits_{\forall\hat{x}^{\prime}\in\hat{X}}\hat{T}(\cdot|\hat{x}_{k},\hat{u},\cdot)V_{0}(\hat{x}^{\prime})\}$;
21    $V_{1}^{\prime}(\hat{x}_{k}):=\max\limits_{\hat{w}\in\hat{W}}{\min\limits_{\hat{A},\hat{T}}}\{\hat{A}(\hat{x}_{k},\hat{u},\cdot)+\sum\limits_{\forall\hat{x}^{\prime}\in\hat{X}}\hat{T}(\cdot|\hat{x}_{k},\hat{u},\cdot)V_{1}(\hat{x}^{\prime})\}$;
22
23   end for
24 $V_{0}:=V_{0}^{\prime},~V_{1}:=V_{1}^{\prime}$;
25
26 end while
[…]
Output: controller $\mathcal{C}=(\mathcal{S},\pi,\mathds{P}_{\psi_{\min}},\mathds{P}_{\psi_{\max}})$
Algorithm 5 Parallel (Pessimistic) Safety Controller Synthesis
Input: IMDP $\hat{\Sigma}$, stopping condition $\varepsilon$
1 $V_{0}:=\textbf{0}_{n}$, $V_{1}:=\textbf{1}_{n}$, $\pi:=\textbf{0}_{n}$;
2 while $\lVert V_{1}-V_{0}\rVert{{}_{\infty}}>\varepsilon$ do
3 for $x\in\mathcal{S}$ in parallel do
4    Solve min LP (5.1) for $\hat{R}$ and $\hat{T}$;
5    $\pi(\hat{x}_{k}):=\operatorname*{argmax}\limits_{\hat{u}\in\hat{U}}\min\limits_{\hat{w}\in\hat{W}}{\min\limits_{\hat{R},\hat{T}}}\{\hat{R}(\hat{x}_{k},\cdot,\cdot)+\sum\limits_{\forall\hat{x}^{\prime}\in\hat{X}}\hat{T}(\cdot|\hat{x}_{k},\cdot,\cdot)V_{0}(\hat{x}^{\prime})\}$;
6    $V_{0}^{\prime}(\hat{x}_{k}):=\max\limits_{\hat{u}\in\hat{U}}\min\limits_{\hat{w}\in\hat{W}}{\min\limits_{\hat{R},\hat{T}}}\{\hat{R}(\hat{x}_{k},\cdot,\cdot)+\sum\limits_{\forall\hat{x}^{\prime}\in\hat{X}}\hat{T}(\cdot|\hat{x}_{k},\cdot,\cdot)V_{0}(\hat{x}^{\prime})\}$;
7    $V_{1}^{\prime}(\hat{x}_{k}):=\max\limits_{\hat{u}\in\hat{U}}\min\limits_{\hat{w}\in\hat{W}}{\min\limits_{\hat{R},\hat{T}}}\{\hat{R}(\hat{x}_{k},\cdot,\cdot)+\sum\limits_{\forall\hat{x}^{\prime}\in\hat{X}}\hat{T}(\cdot|\hat{x}_{k},\cdot,\cdot)V_{1}(\hat{x}^{\prime})\}$;
8
9   end for
10 $V_{0}:=V_{0}^{\prime},~V_{1}:=V_{1}^{\prime}$;
11
12 end while
[…]
14 $V_{0}:=\textbf{0}_{n}$, $V_{1}:=\textbf{1}_{n}$;
15 while $\lVert V_{1}-V_{0}\rVert{{}_{\infty}}>\varepsilon$ do
16 for $x\in\mathcal{S}$ in parallel do
17    Solve max LP (5.1) for $\hat{R}$ and $\hat{T}$;
18      fix $\hat{u}$ from $\pi(\hat{x}_{k})$;
19    $V_{0}^{\prime}(\hat{x}_{k}):=\min\limits_{\hat{w}\in\hat{W}}{\max\limits_{\hat{R},\hat{T}}}\{\hat{R}(\hat{x}_{k},\hat{u},\cdot)+\sum\limits_{\forall\hat{x}^{\prime}\in\hat{X}}\hat{T}(\cdot|\hat{x}_{k},\hat{u},\cdot)V_{0}(\hat{x}^{\prime})\}$;
20    $V_{1}^{\prime}(\hat{x}_{k}):=\min\limits_{\hat{w}\in\hat{W}}{\max\limits_{\hat{R},\hat{T}}}\{\hat{R}(\hat{x}_{k},\hat{u},\cdot)+\sum\limits_{\forall\hat{x}^{\prime}\in\hat{X}}\hat{T}(\cdot|\hat{x}_{k},\hat{u},\cdot)V_{1}(\hat{x}^{\prime})\}$;
21
22   end for
23 $V_{0}:=V_{0}^{\prime},~V_{1}:=V_{1}^{\prime}$;
24
25 end while
[…]
Output: controller $\mathcal{C}=(\mathcal{S},\pi,\mathds{P}_{\psi_{\min}},\mathds{P}_{\psi_{\max}})$
Algorithm 6 Parallel (Pessimistic) Reach-Avoid Controller Synthesis
6.1. Implementations
The performance of the safety, reachability, and reach-avoid specifications can be enhanced by running in parallel all of the nested for loops from Algorithms 3 and 4. These can be seen in the updated Algorithm 5 in Step 333, and Step 141414 for safety specifications. Once more, it is crucial to highlight that we aim to synthesize the minimal $\hat{u}$ as we seek the input that minimizes the likelihood of reaching the region to avoid (i.e., complement of the safe set). Algorithm 6 presents an updated parallel implementation of Algorithm 4 for reach-while-avoid specifications. In particular, Step 3 and Step 14 replace the previous for loops in parallel. We reiterate that when $\mathcal{A}=\emptyset$, the reach-avoid algorithm is reduced to a simpler reachability algorithm.
IMPaCT automatically detects whether to use reachability or reach-avoid specification algorithms based on whether an avoid region has been set by the user. Additionally a Boolean value, IMDP_lower, is set as true for a pessimistic policy, or false for an optimistic policy. For finite-horizon controllers, an additional parameter is needed for the time horizon.
1void setStoppingCondition(double eps);
2void infiniteHorizonReachController(bool IMDP_lower);
3void infiniteHorizonSafeController(bool IMDP_lower);
4void finiteHorizonReachController(bool IMDP_lower, size_t timeHorizon);
5void finiteHorizonSafeController(bool IMDP_lower, size_t timeHorizon);
Listing 10: Synthesis Algorithms.
6.2. Verification Analysis via Interval Markov Chains (IMCs)
In striving for maximum user-friendliness, the tool also automatically distinguishes between verification tasks using an IMC and controller synthesis via an IMDP. In particular, for verification analysis, users only need to specify the absence of input parameters, denoted by dim_u=0~=0= 0. Accordingly, functions like infiniteHorizonReachController, etc., will identify the lack of input and generate a lookup table represented as $\mathcal{C}:=(\mathcal{S},\mathds{P}_{\psi_{\min}},\mathds{P}_{\psi_{\max}})$, although it will still be saved as controller.h5. In fact, the verification’s lookup table provides both the minimum and maximum probabilities of fulfilling a desired property corresponding to each partition of the finite system.
7. Loading and Saving Files
We use HDF5 [The23] as the data format for saving and loading into IMPaCT. In particular, HDF5 is a common widely supported format for large, heterogeneous, and complex data sets. This data structure is self-descriptive, eliminating the need for extra metadata to interpret the files. Additionally, it supports “data slicing”, enabling extraction of specific segments from a dataset without the necessity of analyzing the entire set.
The primary advantage of HDF5 lies in its open format, which ensures native support across numerous programming languages and tools, such as MATLAB, Python, and R. This should facilitate simpler sharing of synthesized controllers without requiring end-users to install additional programs. In IMPaCT, the subsequent commands are employed to save data in HDF5 format:
1void saveStateSpace();
2void saveInputSpace();
3void saveDisturbSpace();
4void saveTargetSpace();
5void saveAvoidSpace();
6void saveMinTargetTransitionVector();
7void saveMaxTargetTransitionVector();
8void saveMinAvoidTransitionVector();
9void saveMaxAvoidTransitionVector();
10void saveMinTransitionMatrix();
11void saveMaxTransitionMatrix();
12void saveController();
Listing 11: Save functions.
Furthermore, users have the flexibility to build the abstraction using their preferred methods and load the required matrices into IMPaCT for verification or controller synthesis tasks. For an illustration, we refer to the example ex_load_reach.
1void loadStateSpace(string filename);
2void loadInputSpace(string filename);
3void loadDisturbSpace(string filename);
4void loadTargetSpace(string filename);
5void loadAvoidSpace(string filename);
6void loadMinTargetTransitionVectorx(string filename);
7void loadMaxTargetTransitionVector(string filename);
8void loadMinAvoidTransitionVector(string filename);
9void loadMaxAvoidTransitionVector(string filename);
10void loadMinTransitionMatrix(string filename);
11void loadMaxTransitionMatrix(string filename);
12void loadController(string filename);
Listing 12: Load functions.
8. Benchmarking and Case Studies
We illustrate IMPaCT’s applications by employing multiple well-known benchmark systems, encompassing safety, reachability, and reach-avoidance specifications across infinite horizons. Among these, we leverage several complex physical case studies, encompassing a 2D robot, a 3D autonomous vehicle, 3D and 5D room temperature control systems, as well as 4D and 7D building automation systems. We also consider a 14D case study used for the purposes of showing the scalability of the tool. All the case studies are run on a high performance computer with 4 Intel Xeon E7-8857 v2 CPUs @ 3.6GHz. In Table 1, we showcase the computational times for these case studies, all executed over an infinite time horizon.
Remark 8.1.
It is worth mentioning that the parallelization in IMPaCT relies on external function libraries for optimization and dynamic programming. Consequently, the GPU parallelization, following our proposed algorithms, is currently in progress and will soon be integrated into the Git repository.
8.1. 2D Robot
Consider a robot described by the following nonlinear difference equations:
$\begin{bmatrix}x_{1}(k+1)\\ x_{2}(k+1)\end{bmatrix}=\begin{bmatrix}x_{1}(k)+10u_{1}(k)\cos(u_{2}(k))+w(k)+\varsigma_{1}(k)\\ x_{2}(k)+10u_{2}(k)\sin(u_{2}(k))+w(k)+\varsigma_{2}(k)\end{bmatrix}\!\!,$
where $(x_{1},x_{2})\in X:=[-10,10]^{2}$ is the spacial coordinate of the location of the robot, $(u_{1},u_{2})\in U:=[-1,1]^{2}$ is the input, and $w\in W:=[-0.5,0.5]$ is the disturbance. The noise $(\varsigma_{1},\varsigma_{2})$ has the covariance matrix ${\mathsf{Cov}}:=\mathsf{diag}(0.75,0.75)$. The discrete intervals are, respectively, $\eta_{x}=(0.5,0.5)$, $\eta_{u}=(0.1,0.1)$ and $\eta_{w}=0.1$. The target set $\mathcal{T}:=[5,7]^{2}$ and the avoid set $\mathcal{A}:=[-2,2]^{2}$.
We consider four scenarios for this case study:
  • Reach-avoid specification with no disturbance, see Fig. 1(B), i.e., $\hat{w}(k)=0$.
  • Reach-avoid specification with disturbance.
  • Reachability specification with no disturbance with updated $\eta_{x}=(1,1)$ and $\eta_{u}=(0.2,0.2)$, see Fig. 1(A).
  • Reachability specification with disturbance with with updated $\eta_{x}=(1,1)$, $\eta_{u}=(0.2,0.2)$.
The latter two scenarios are designed for swift simulations on small personal computers.
(a) 2D Robot - Reachability
(b) 2D Robot - Reach-Avoid
Figure 1. 2D Robot case study fulfilling reachability and reach-avoid properties with different noise realizations. The green and red boxes are target and avoid regions, respectively.
8.2. 3D Autonomous Vehicle
Consider a 3-dimensional autonomous vehicle described by the following nonlinear difference equations:
$\begin{bmatrix}x_{1}(k+1)\\ x_{2}(k+1)\\ x_{3}(k+1)\end{bmatrix}=\begin{bmatrix}x_{1}(k)+(u_{1}(k)\cos(\alpha+x_{3}(k))\cos(\alpha)^{-1})T_{s}+{\varsigma_{1}(k)}\\ x_{2}(k)+(u_{1}(k)\sin(\alpha+x_{3}(k))\cos(\alpha)^{-1})T_{s}+{\varsigma_{2}(k)}\\ x_{3}(k)+(u_{1}(k)\tan(u_{2}(k)))T_{s}+{\varsigma_{3}(k)}\end{bmatrix}\!\!,$
where $\alpha=\arctan(\frac{\tan(u_{2})}{2})$, and $T_{s}=0.1$ is the sampling time. The inputs $(u_{1},u_{2})\in U:=[-1,4]\times[-0.4,0.4]$ represent the wheel velocity and the steering angle, where $\eta_{u}=(1,0.2)$. The states $(x_{1},x_{2})$ are the spatial coordinates, and $x_{3}$ is the orientation, where $(x_{1},x_{2},x_{3})\in X:=[-5,5]^{2}\times[-3.4,3.4]$ and $\eta_{x}=(0.5,0.5,0.4)$. The noise $(\varsigma_{1},\varsigma_{2},\varsigma_{3})$ has covariance matrix ${\mathsf{Cov}}:=\mathsf{diag}(\frac{2}{3},\frac{2}{3},\frac{2}{3})$. The target set $\mathcal{T}$ and avoid set $\mathcal{A}$ are described by the hyper-rectangles [5.75,0.25]×[0.25,5.75]×[3.45,3.45][-5.75,0.25$\times$[-0.25,5.75$\times$[-3.45,3.45][ - 5.75 , 0.25 ] × [ - 0.25 , 5.75 ] × [ - 3.45 , 3.45 ] and [5.75,0.25]×[0.75,0.25]×[3.45,3.45][-5.75,0.25$\times$[-0.75,-0.25$\times$[-3.45,3.45][ - 5.75 , 0.25 ] × [ - 0.75 , - 0.25 ] × [ - 3.45 , 3.45 ], respectively. For this case study, we consider a reach-while-avoid specification with no disturbances, see Fig. 2.
Figure 2. 3D Autonomous Vehicle case study fulfilling reach-while-avoid property, with different noise realizations, starting from an initial condition [3;3;0.6][3;-3;0.6][ 3 ; - 3 ; 0.6 ]. The green and red boxes are target and avoid regions, respectively.
8.3. 3D and 5D Room Temperature Control Systems
We also employ IMPaCT for the room temperature control systems taken from the ARCH competition [ABC+], one with 333 dimensions and another one with 555 dimensions. The two case studies both consider two inputs described by $(u_{1},u_{2})\in U:=[0,1]^{2}$ with $\eta_{u}=(0.2,0.2)$, and have a state space described by $(x_{1},\ldots,x_{n})\in X:=[19,21]^{n}$, where $n$ is the state dimension. For the  [^3] D case, the dynamics are described as:
$\begin{bmatrix}x_{1}(k+1)\\ x_{2}(k+1)\\ x_{3}(k+1)\end{bmatrix}\!=\!\begin{bmatrix}(a-\gamma u_{1}(k))x_{1}(k)+{\zeta}(x_{2}(k)+x_{3}(k))+\gamma T_{h}u_{1}(k)+\beta T_{e}+\varsigma_{1}(k)\\ ax_{2}(k)+{\zeta}(x_{1}(k)+x_{3}(k))+\beta T_{e}+\varsigma_{2}(k)\\ (a-\gamma u_{2}(k))x_{3}(k)+{\zeta}(x_{1}(k)+x_{2}(k))+\gamma T_{h}u_{2}(k)+\beta T_{e}+\varsigma_{3}(k)\end{bmatrix}\!\!,$
where $\beta=0.022,\gamma=0.05,{\zeta}=0.2$, are the conduction factors, respectively, between the external environment and the current room, between the heater and the current room, and between neighbouring rooms. In addition, $T_{h}=50^{\circ}C$ is the heater temperature, $T_{e}=-1^{\circ}C$ is the outside temperature, and $a=1-2{\zeta}-\beta$. Furthermore, $\eta_{x}=(0.1,0.1,0.1)$ and the noise has covariance matrix ${\mathsf{Cov}}:=\mathsf{diag}(0.02,0.02,0.02)$.
Figure 3. 3D Room Temperature fulfilling safety properties over 101010 time steps, with different noise realizations, starting from an initial condition [21;21;21][21;21;21][ 21 ; 21 ; 21 ].
The dynamics for the  [^5] D case study are characterized as:
$\begin{bmatrix}x_{1}(k+1)\\ x_{2}(k+1)\\ x_{3}(k+1)\\ x_{4}(k+1)\\ x_{5}(k+1)\end{bmatrix}\!=\!\begin{bmatrix}(a-\gamma u_{1}(k))x_{1}(k)+{\zeta}(x_{2}(k)+x_{5}(k))+\gamma T_{h}u_{1}(k)+\beta T_{e}+\varsigma_{1}(k)\\ ax_{2}(k)+{\zeta}(x_{1}(k)+x_{3}(k))+\beta T_{e}+\varsigma_{2}(k)\\ (a-\gamma u_{2}(k))x_{3}(k)+{\zeta}(x_{4}(k)+x_{2}(k))+\gamma T_{h}u_{2}(k)+\beta T_{e}+\varsigma_{3}(k)\\ ax_{4}(k)+{\zeta}(x_{3}(k)+x_{5}(k))+\beta T_{e}+{\varsigma_{4}(k)}\\ ax_{5}(k)+{\zeta}(x_{4}(k)+x_{1}(k))+\beta T_{e}+{\varsigma_{5}(k)}\end{bmatrix}\!\!,$
where ${\zeta}=0.3$ and $\eta_{x}=(0.4,0.4,0.4,0.4,0.4)$. The noise has covariance matrix ${\mathsf{Cov}}:=\mathsf{diag}(0.01,0.01,0.01,0.01,0.01)$.
Both case studies examine safety specifications where the safe region $\mathcal{S}=X$. Figure 3 showcases a simulation of the safety controller spanning 101010 time steps, while Table 1 presents the synthesis times for this case study within an infinite time horizon.
8.4. 4D and 7D Building Automation Systems
The following benchmarks are based on a smart building at the University of Oxford [CA18]. We consider two versions of this benchmark with different dimensions. The first one involves a safety synthesis problem rooted in a two-zone configuration, featuring stochastic dynamics:
$x(k+1)=\begin{bmatrix}0.6682&0&0.02632&0\\ 0&0.683&0&0.02096\\ 1.0005&0&-0.000499&0\\ 0&0.8004&0&0.1996\end{bmatrix}x(k)+\begin{bmatrix}0.1320\\ 0.1402\\ 0\\ 0\end{bmatrix}u(k)+\begin{bmatrix}3.4378\\ 2.9272\\ 13.0207\\ 10.4166\end{bmatrix}+\varsigma(k),$
where the covariance matrix ${\mathsf{Cov}}:={\mathsf{diag}}(12.9199,12.9199,2.5826,3.2279)$ and the off-diagonal elements are zero. Moreover, $x\in X:=[19,21]^{2}\times[30,36]^{2}$ with $\eta_{x}:=(0.5,0.5,1.0,1.0)$, and $u\in U:=[17,20]$ with $\eta_{u}:=(1.0,1.0)$, see Fig. 4.
Figure 4. 4D Building Automation System fulfilling safety properties within 101010 time steps, with 555 different noise realizations, starting from an initial condition [19;20;36;36][19;20;36;36][ 19 ; 20 ; 36 ; 36 ].
The second version is a safety verification problem (i.e., without input variables), in which the dynamics extend the first benchmark to consider a larger number of continuous variables:
$x(k+1)=\begin{bmatrix}0.968&0&0.004&0&0.004&0&0.004\\ 0&0.968&0&0.003&0&0.003&0.003\\ 0.011&0&0.949&0&0&0&0\\ 0&0.010&0&0.952&0&0&0\\ 0.011&0&0&0&0.949&0&0\\ 0&0.010&0&0&0&0.952&0\\ 0.011&0.010&0&0&0&0&0.979\end{bmatrix}x(k)+\varsigma(k),$
where the noise has covariance ${\mathsf{Cov}}:={\mathsf{diag}}(51.3,50.0,21.8,23.5,25.2,26.5,91.7)$, with off-diagonals equal to zero. In addition, $x(k)\in X:=[-0.5,0.5]^{7}$ with $\eta_{x}:=(0.05,0.05,0.5,0.5,0.5,0.5,0.5)$.
It is noteworthy that, as indicated in Table 1, the construction of the IMDP abstraction for the  [^7] D building automation systems was accomplished within 242424 hours. However, the synthesis process extended beyond 242424 hours to complete. This prolonged duration is primarily attributed to the expansive size of the state space, comprising 107,163107,163107 , 163 finite states, while each state functions as a decision variable within the dynamic program. Considering that the synthesis involves solving the dynamic program for each row, the number of decision variables in one iteration amounts to $107,163^{2}$, i.e., approximately 11.511.511.5 billion decision variables.
8.5. 14D Case Study
To demonstrate the scalability of IMPaCT, we consider a 14D case study, borrowed from the relevant literature [CA19, LKSZ20], with the following dynamics:
$x(k+1)=0.8x(k)+0.2\varsigma(k),$ (8.1)
where $x=[x_{1};\dots;x_{14}],\varsigma=[\varsigma_{1};\dots;\varsigma_{14}]$, with $X:=[-0.5,0.5]^{14}$. We consider this case study to solve a safety verification problem, as detailed in Table 1.
8.6. Comparison with StocHy
As previously highlighted in the introduction, StocHy utilizes the value iteration algorithm for IMDP construction, which lacks convergence guarantees when dealing with infinite-horizon specifications. In contrast, IMPaCT employs the interval iteration algorithm to guarantee convergence to an optimal controller. This crucial feature, in addition to offering parallelization, distinguishes IMPaCT from StocHy, which defaults to value iteration without general parallelization capabilities.
Beyond this guarantee, we aim to execute the  [^14] D case study using StocHy to showcase the efficiency of our tool. While IMPaCT completed the IMDP construction within an hour (535353 minutes), as detailed in Table 1, StocHy failed to complete task within a 242424-hour timeframe.
9. Conclusion
In this work, we developed the advanced software tool IMPaCT, which is the first tool to exclusively construct IMC/IMDP abstraction and perform verification and controller synthesis over infinite-horizon properties while providing convergence guarantees. Developed in C++ using AdaptiveCpp, an independent open-source implementation of SYCL, IMPaCT capitalizes on adaptive parallelism across diverse CPUs/GPUs of all hardware vendors, including Intel and NVIDIA. We have benchmarked IMPaCT across various physical case studies including a 2D robot, a 3D autonomous vehicle, a 5D room temperature control system, and a 7D building automation system, borrowed from the ARCH tool competition, with its scalability further highlighted through a 14D case study.
Acknowledgment
The authors express their gratitude to Sadegh Soudjani for valuable discussions at the early stage of this project.